Determine the [OH−] of a 2.3x10^−3 M solution of HNO3 ​in units x10^-12
Answer #1

HNO3 is a strong salt

Thus, [HNO3] = [H+] = 2.3 *10-3 M

Now, [H+][OH-] = 10-14

(2.3 *10-3M) [OH-] = 10-14

[OH-] = 4.35 *10-12 M